OpenStudy (kate_kat):

PLZ HELP! sorry if this is a confusing question, if you can answer it right i will medal! What is the equation of this line? a line that has points at (-8,2) and (8,-2)

OpenStudy (welshfella):

given the 2 poits we can find the slope of the line joining the points. Then we can sue the point-slope general form of the equation of a straight line y - y1 = m(x - x1) where m = slope and (x1,y1) is any point on the line.

OpenStudy (kate_kat):

answers are... A.y=-4x B.y=4x C. y=-1/4x D. 1/4x, im thinking its A

OpenStudy (welshfella):

* we can use

OpenStudy (welshfella):

first work out the slope which can be done using the pythagoras theorem